OpenStudy (anonymous):

P(-2,-1) m=2 write an equation of the line perpendicular to y=2x+3 passing through the origin

OpenStudy (anonymous):

What is the given point P(-2,-1) .? Is that given in th problem?

OpenStudy (anonymous):

yes that given point is p(-2,-1)

OpenStudy (anonymous):

You are asked to write equation perpendicular to y=2x+3 passing thru origin, right?

OpenStudy (anonymous):

In y = mx+b, b is the y-intercept. So since it will pass thru the origin b=0

OpenStudy (anonymous):

When 2 lines are perpendicular, their slopes. Are negative reciprocals. Since slope of the given line is 2, what is the negative reciprocal of 2?

OpenStudy (anonymous):

1/2 right

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

You can now write the equation

OpenStudy (anonymous):

m = -1/2 and b =0

OpenStudy (anonymous):

so : y=-1/2 x

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

thank you!(: